OpenStudy (anonymous):

If (goh)(x)=x^2+6x+2 and h(x)=x+1, what is g(x)?

OpenStudy (anonymous):

@terenzreignz mali given nia, xDDD

terenzreignz (terenzreignz):

But I thought this was already sorted out? :( Oh well...

terenzreignz (terenzreignz):

You know that \[\Large (g\circ h)(x) = g(h(x))\] right?

terenzreignz (terenzreignz):

binibitin mo na naman ako...-.-

OpenStudy (anonymous):

ahah, xD mali 4x pala talga un. xDD

OpenStudy (anonymous):

di ko maintindihan sagot nung isa

OpenStudy (anonymous):

If (goh)(x)=x^2+4x+2 and h(x)=x+1, what is g(x)? tama pala yan